# for Unseen Language Generalization

### 11 Jan 2024 v1

@@ Abstract
Pretrained language models (PLMs) have shown remarkable generalization toward multiple tasks and languages. Nonetheless, the generalization of PLMs towards unseen languages is poor, resulting in significantly worse language performance, or even generating nonsensical responses that are comparable to a random baseline. This limitation has been a long-standing problem of PLMs raising the problem of diversity and equal access to language modeling technology. In this work, we solve this limitation by introducing LinguAlchemy, a regularization technique that incorporates various aspects of languages covering typological, geographical, and phylogenetic constraining the resulting representation of PLMs to better characterize the corresponding linguistics constraints. LinguAlchemy significantly improves the accuracy performance of mBERT and XLM-R on unseen languages by $\sim$18and $\sim$2%, respectively compared to fully fine-tuned models and displaying a high degree of unseen language generalization. We further introduce AlchemyScale and AlchemyTune, extension of LinguAlchemy which adjusts the linguistic regularization weights automatically, alleviating the need for hyperparameter search. LinguAlchemy enables better cross-lingual generalization to unseen languages which is vital for better inclusivity and accessibility of PLMs.
@@ 1 Introduction
To overcome those limitations, in this work, we introduce LinguAlchemy. Unlike adapter-based approaches which isolate the capability to understand different languages over multiple language-specific adapters, LinguAlchemy learns a shared representation across multiple languages, allowing the model to leverage shared knowledge between languages. Such behavior is attainable by eliminating the language-specific module from the model and instead utilizing a regularization to learn language-specific knowledge. LinguAlchemy allows the model to perform inference without knowing the language of the query prior. Our evaluation suggests that LinguAlchemy can improve XLM-R generalization in unseen languages while maintaining the performance on the high-resource languages without requiring information on the language of the query apriori.
@@ 2 Related Works
PLMs with their transformer-based architectures have been demonstrating exceptional capabilities in language comprehension and generation. These models excel in abstract linguistic generalization by capturing complex linguistic patterns and understanding structural positions and thematic roles, which are crucial for interpreting language semantics. Research in this area Ganesh et al. (2021) has provided critical insights that enable these models to process and generate human language effectively. The studies have explored how these models grasp intricate linguistic features, including syntax and semantics, thereby enhancing their performance across a wide range of language tasks .
@@
However, despite these advancements, PLMs still face significant challenges in generalizing to unseen languages, particularly when adapting to low-resource and unseen languages. These challenges stem from the vast structural and semantic variation across languages Bender ; Jurafsky and Martin (2019), the scarcity of resources Mohammad (2019); Lewis et al. (2020), and the limitations inherent in the models themselves Lin et al. (2017); Vaswani et al. (2017). This situation highlights the complexity of scaling and generalizing these models effectively and underscores the need for more sophisticated approaches in model training and adaptation to ensure broader and more equitable language coverage.
@@ 3.1 Does Multilingual LMs capture Linguistic Constraint?
These feature vectors denote a typological feature that is adapted from several sources including World Atlas of Language Structures (WALS), Syntactic Structures of World Languages Collins and Kayne (2009), and short prose descriptions on typological features in Ethnologue Lewis (2009). Syntax vectors captures information about the syntactic properties of languages which derived from large-scale typological databases, which document the structural and semantic variation across different languages. These syntax features in the URIEL vector are utilized to represent languages in vector form, allowing for the analysis and comparison of languages based on their syntactic properties.
@@ 3.2 Proof of Concept
Algorithm 1 Language Representation and Alignment Process
0:  Aligned Language Representations $A$
  $F_{U}\leftarrow$ ExtractFeatures($U$)
  $S\leftarrow\{\}$
  for each sentence $x$ in $D$ do
     $H_{x}\leftarrow$ GetLastHiddenStates($x$, $\Theta$)
     $N_{x}\leftarrow$ CountAttentionMasked($x$)
     $R_{x}\leftarrow\frac{\textsc{Sum}(H_{x})}{N_{x}}$
     $S\leftarrow S\cup\{R_{x}\}$
  end for
  $W,b\leftarrow$ InitializeParameters()
  for each training epoch do
     $P_{U}\leftarrow(W\times S)+b$
     $loss\leftarrow$ ComputeLoss($P_{U}$, $F_{U}$)
     $W,b\leftarrow$ UpdateParametersWithConstraint($W,b,loss$)
  end for
  $A\leftarrow\{\}$
  for each sentence representation $s$ in $S$ do
     $A_{m}\leftarrow(W\times s)+b$
     $A\leftarrow A\cup\{A_{m}\}$
  end for
Figure 1 presents a visual analysis facilitated by UMAP McInnes et al. (2018), showing the correlation between mBERT language representation and the linguistic vectors from the URIEL database ($R^{2}$ = 0.816). By leveraging UMAP, the plot accentuates the principal variances within the joint feature space of the embeddings and vectors. The spatial representation of languages on this plot mirrors their linguistic and geographical relatedness, as encapsulated by mBERT. This visualization underscores the model’s ability to mirror linguistic typologies, with languages sharing common roots such as ’de-DE’ and ’nl-NL’ naturally clustering together. The density and arrangement of these clusters potentially reflect mBERT capacity to capture and represent language family traits. Conversely, the presence of sparser clusters or outliers prompts a closer examination of mBERT’s coverage and consistency in representing diverse linguistic features.
@@
where $R$ represents the model-generated representations, $U$ denotes the URIEL vectors, and $N$ is the number of data points.
@@ 5 Results and Discussion
mBERT XLM-R
Language AF Full FT Ours Full FT Ours
am-ET* 4.61% 5.45% 58.07% 78.55% 77.00%
cy-GB 25.05% 23.81% 30.01% 64.39% 68.96%
af-ZA 57.67% 52.72% 50.16% 82.72% 75.73%
km-KH* 7.83% 8.34% 59.85% 84.57% 78.70%
sw-KE 22.16% 19.77% 54.93% 58.07% 74.90%
mn-MN* 27.57% 27.17% 57.38% 87.53% 76.27%
tl-PH 40.25% 37.46% 66.47% 85.91% 80.42%
kn-IN 40.99% 34.23% 67.84% 80.46% 81.19%
te-IN 34.40% 35.27% 71.88% 84.57% 82.59%
sq-AL 49.53% 44.75% 70.66% 67.92% 82.17%
ur-PK 47.11% 42.80% 69.43% 73.60% 81.81%
az-AZ 63.82% 61.57% 69.18% 80.16% 82.03%
ml-IN 35.84% 27.71% 67.83% 78.92% 81.80%
ms-MY 65.80% 66.48% 67.85% 82.99% 81.80%
ca-ES 73.07% 73.10% 68.37% 87.36% 82.04%
sl-SL 49.29% 47.18% 68.52% 86.28% 82.15%
sv-SE 64.09% 60.05% 68.36% 85.37% 82.20%
ta-IN 41.66% 34.87% 68.63% 84.43% 82.38%
nl-NL 69.97% 70.65% 68.62% 81.98% 82.34%
it-IT 71.86% 70.78% 68.14% 78.31% 82.14%
he-IL 51.24% 48.18% 68.07% 88.74% 82.25%
pl-PL 62.27% 59.99% 67.05% 61.30% 81.60%
da-DK 71.25% 71.72% 66.39% 76.46% 81.43%
nb-NO 68.76% 68.53% 65.71% 78.18% 81.33%
ro-RO 58.68% 54.20% 64.93% 82.82% 81.34%
th-TH 30.43% 24.21% 64.37% 73.30% 81.05%
fa-IR 59.35% 56.89% 64.38% 77.20% 81.04%
Average 47.95% 45.48% 64.19% 78.97% 80.32%
@@ 5.1 LinguAlchemy Performance
To evaluate the effectiveness of our proposed technique on unseen languages, we trained mBERT and XLM-R on the MASSIVE dataset. Specifically, we train on 25 languages and test on 27 different, unseen languages. Our results are summarized in Table 1. We compared our method with zero-shot generalization, where the model is fully tuned on seen languages and then tested on unseen languages (referred to as Full FT in the Table). Additionally, we explored AdapterFusion Pfeiffer et al. (2021a) as another baseline (AF in the Table). AdapterFusion has shown better adaptation to unseen languages than naive zero-shot generalization. Unfortunately, many language adapters that we need for AdapterFusion is not available for XLM-R.
@@
Figure 2: Average performance of unseen languages under dynamic URIEL loss scales
@@ 5.2 Effect of Scaling URIEL loss
One issue with introducing a scaling factor is the addition of another tunable hyperparameter. Intuitively, we might aim for a balanced weight between the classification and URIEL losses. Therefore, instead of expensively testing different scaling factors, an adaptive scaling factor might be more cost-effective and beneficial.
Here, we explore two ideas: dynamic and trainable factors. The results of these approaches can be seen in Table 2.
@@
Dynamic 62.97% 80.43%
Trainable 63.24% 79.10%
Table 2: Average performance of unseen languages under dynamic URIEL loss scaling factors
@@ 6 Conclusion
Despite the promising results achieved by LinguAlchemy in enhancing the performance across a wide range of unseen languages in intent classification tasks, there are notable limitations in our approach. Firstly, the performance on seen languages remains suboptimal, suggesting a potential area of improvement, possibly through methods such as weight freezing. Secondly, we observe a trade-off where enhancing generalization capabilities seems to come at the cost of reduced accuracy in seen languages. This warrants further investigation into methodologies that can balance performance across both seen and unseen languages more effectively. Additionally, our research is currently limited to intent classification tasks, and extending this approach to a broader range of NLP tasks could provide deeper insights into its versatility and applicability. Another limitation lies in the selection of URIEL features based on theoretical reasoning; while our choice of syntax, geography, and language family features is grounded in Chapter 3, empirical testing with alternative or additional features could uncover opportunities for further refinement of the model. Addressing these limitations could significantly enhance the generalizability and overall performance of multilingual models in NLP applications.
@@ Appendix A Languages in Dataset
The MASSIVE Dataset or Multilingual Amazon SLU Resource Package (SLUPR) which encompasses slot-filling, intent classification, and Virtual Assistant Evaluation tasks, is a new dataset containing about 1 million realistic data examples shaped from utterances by annotators. This dataset spans 51 languages, 60 intents, 55 slot types, and 18 domains.

### 01 Feb 2024 v2

@@ Abstract
Pretrained language models (PLMs) have become remarkably adept at task and language generalization. Nonetheless, they often fail dramatically when faced with unseen languages, posing a significant problem for diversity and equal access to PLM technology. In this work, we present LinguAlchemy, a regularization technique that incorporates various aspects of languages covering typological, geographical, and phylogenetic constraining the resulting representation of PLMs to better characterize the corresponding linguistics constraints. LinguAlchemy significantly improves the accuracy performance of mBERT and XLM-R on unseen languages by $\sim$18and $\sim$2%, respectively compared to fully fine-tuned models and displaying a high degree of unseen language generalization. We further introduce AlchemyScale and AlchemyTune, extension of LinguAlchemy which adjusts the linguistic regularization weights automatically, alleviating the need for hyperparameter search. LinguAlchemy enables better cross-lingual generalization to unseen languages which is vital for better inclusivity and accessibility of PLMs.
@@ 1 Introduction
Figure 1: LinguAlchemy enhances performance in unseen languages by allowing the model to predict the linguistic vector and then fitting it via a similarity loss towards the specific language’s URIEL vector.
@@
To overcome those limitations, in this work, we introduce LinguAlchemy. Unlike adapter-based approaches which isolate the capability to understand different languages over multiple language-specific adapters, LinguAlchemy learns a shared representation across multiple languages, allowing the model to leverage shared knowledge between languages. Such behavior is attainable by eliminating the language-specific module from the model and instead utilizing a regularization to learn language-specific knowledge. LinguAlchemy allows the model to perform inference without knowing the language of the query prior. Our evaluation suggests that LinguAlchemy can improve mBERT and XLM-R generalization in unseen languages while maintaining the performance on the high-resource languages without requiring information on the language of the query apriori.
@@ 2 Related Works
PLMs with their transformer-based architectures have been demonstrating exceptional capabilities in language comprehension and generation. These models excel in abstract linguistic generalization by capturing complex linguistic patterns and understanding structural positions and thematic roles, which are crucial for interpreting language semantics. Research in this area Ganesh et al. (2021) has provided critical insights that enable these models to process and generate human language effectively. The studies have explored how these models grasp intricate linguistic features, including syntax and semantics, thereby enhancing their performance across a wide range of language tasks Rathore et al. (2023).
@@
However, despite these advancements, PLMs still face significant challenges in generalizing to unseen languages, particularly when adapting to low-resource and unseen languages. These challenges stem from the vast structural and semantic variation across languages Bender (2011); Jurafsky and Martin (2019), the scarcity of resources Mohammad (2019); Lewis et al. (2020), and the limitations inherent in the models themselves Lin et al. (2017). This situation highlights the complexity of scaling and generalizing these models effectively and underscores the need for more sophisticated approaches in model training and adaptation to ensure broader and more equitable language coverage.
@@ 3.1 Does Multilingual LMs capture Linguistic Constraint?
These feature vectors denote a typological feature that is adapted from several sources including World Atlas of Language Structures (WALS), Syntactic Structures of World Languages Collins (2010), and short prose descriptions on typological features in Ethnologue Lewis (2009). Syntax vectors captures information about the syntactic properties of languages which derived from large-scale typological databases, which document the structural and semantic variation across different languages. These syntax features in the URIEL vector are utilized to represent languages in vector form, allowing for the analysis and comparison of languages based on their syntactic properties.
@@ 3.2 Proof of Concept
Figure 2 presents a visual analysis facilitated by UMAP McInnes et al. (2018), showing the correlation between mBERT language representation and the linguistic vectors from the URIEL database ($R^{2}$ = 0.816). By leveraging UMAP, the plot accentuates the principal variances within the joint feature space of the embeddings and vectors. The spatial representation of languages on this plot mirrors their linguistic and geographical relatedness, as encapsulated by mBERT. This visualization underscores the model’s ability to mirror linguistic typologies, with languages sharing common roots such as ’de-DE’ and ’nl-NL’ naturally clustering together. The density and arrangement of these clusters potentially reflect mBERT capacity to capture and represent language family traits. Conversely, the presence of sparser clusters or outliers prompts a closer examination of mBERT’s coverage and consistency in representing diverse linguistic features. We also formally defined the language representation alignment in the Appendix B.
@@
where $R$ represents the model-generated representations, $U$ denotes the URIEL vectors, and $N$ is the number of data points. To generate the model representation, we take the output representation from the $CLS$ token and multiply it with a new, trainable projection layer to transform the vector size so that they are compatible.
@@ 5.1 LinguAlchemy Performance
To evaluate the effectiveness of our proposed technique on unseen languages, we trained mBERT and XLM-R on the MASSIVE dataset. Specifically, we train on 25 languages and test on 27 different, unseen languages. Our results are summarized in Table 1. We compared our method with zero-shot generalization, where the model is fully tuned on seen languages and then tested on unseen languages (referred to as Full FT in the Table). Additionally, we explored AdapterFusion Pfeiffer et al. (2021a) as another baseline. AdapterFusion has shown better adaptation to unseen languages than naive zero-shot generalization. Unfortunately, many language adapters that we need for AdapterFusion is not available for XLM-R.
@@ 5.2 Effect of Scaling URIEL loss
Figure 3: Average performance of unseen languages under various URIEL loss scales.
@@
One issue with introducing a scaling factor is the addition of another tunable hyperparameter. Intuitively, we might aim for a balanced weight between the classification and URIEL losses. Therefore, instead of expensively testing different scaling factors, an adaptive scaling factor might be more cost-effective and beneficial. Here, we explore two ideas: dynamic and trainable factors. The results of these approaches can be seen in Table 2.
@@
AlchemyScale 62.97% 80.43%
AlchemyTune 63.24% 79.10%
Table 2: Performance Comparison Across Different URIEL Scaling Methods.
@@
Train Group Lang. Family Languages Num. Languages
1 Indo-European
af-ZA, bn-BD, ca-ES, cy-GB, da-DK, de-DE, el-GR, en-US, es-ES, fa-IR, fr-FR, hi-IN, hy-AM, is-IS, it-IT, lv-LV, nb-NO, nl-NL, pl-PL, pt-PT, ro-RO, ru-RU, sl-SL, sq-AL, sv-SE, ur-PK
26
2 Dravidian
Train Group 1 + kn-IN, ml-IN, ta-IN, te-IN
30
3 Afro-Asiatic
Train Group 2 + am-ET, ar-SA, he-IL
33
4 Sino-Tibetan
Train Group 3 + my-MM, zh-CN, zh-TW
36
Unseen Languages
sw-KE, km-KH, vi-VN, id-ID, jv-ID, ms-MY, tl-PH, ja-JP, ka-GE, ko-KR, mn-MN, th-TH, az-AZ, tr-TR, fi-FI, hu-HU
16
Table 3: Language family distribution used in the language family generalization experiment (§5.3)
5.3 Generalization Across Language Family
We investigate LinguAlchemy across language families to further analyze the generalization capabilities of BERT and XLM-R models. This experiment offers insight into how adaptable LinguAlchemy is to a variety of linguistic features. We perform our experiment by splitting the languages in MASSIVE according to their language families and train the model on a subset of language families while testing on the rest, unseen language families. We explore on including different subset of language families, as seen in Table 3.
The "others unseen" category includes additional language families not incorporated in the training set, serving as an "unseen" testbed. As illustrated in Figure 4, LinguAlchemy demonstrates generalization towards these unseen language families. Perhaps unsurprisingly, adding more languages and, importantly, diversity to the training data improves generalization performance. Interestingly, including the Afro-Asiatic language family group—comprising "am-ET," "ar-SA," and "he-IL," all of which have distinct scripts—resulted in a notable performance enhancement from train group 2 to train group 3 experiment.
The performance of both models, combined with LinguAlchemy underscores the advantage of including a broader spectrum of languages within training groups for enhanced model generalization. However, the impact of this diversity is not uniform across all language families: While some consistently benefit from the expansion of training data, others do not, indicating that merely increasing the volume of data from the same family may not necessarily improve performance. This inconsistency indicates the potential limitations within the models’ capacity to learn and generalize the linguistic features specific to certain language families. Consequently, our observation shows that the degree of generalization varies noticeably among different families, suggesting that while some may significantly profit from these models’ capabilities, others may require more tailored strategies to gain similar performance improvement.
Figure 4: Model performance across language families. Dotted lines indicates language families used in training in some of the training stages (solid dots for active use–refer to Table 3), and solid grey lines for families unseen in all training stages, with variance shown in shading.
@@ 6 Conclusion
LinguAlchemy enhances performance across many unseen languages in intent classification, yet it faces limitations. Performance on seen languages is less than ideal, indicating room for improvement through methods like weight freezing. Also, better generalization appears to reduce accuracy in seen languages, pointing to a need for balanced approaches. Currently, the research is limited to intent classification, and expanding to other NLP tasks could reveal more about its versatility. Moreover, the choice of URIEL features—syntax, geography, language family—is theoretically sound, as discussed in Chapter 3, but empirical tests with different features might refine the model further. Overcoming these limitations could greatly improve the generalizability and effectiveness of multilingual NLP models.
@@ Appendix A Languages in Dataset
The MASSIVE Dataset, also known as the Multilingual Amazon SLU Resource Package (SLUPR), offers a comprehensive collection of approximately one million annotated utterances for various natural language understanding tasks such as slot-filling, intent detection, and Virtual Assistant performance evaluation. It is an extensive dataset that includes 51 languages, 60 intents, 55 slot types, and spans 18 different domains. The dataset is further enriched with a substantial amount of English seed data, comprising 587k training utterances, 104k development utterances, and 152k test utterances.
@@
Appendix B Algorithm
Algorithm 1 Language Representation and Alignment Process
0:  Aligned Language Representations $A$
  $F_{U}\leftarrow$ ExtractFeatures($U$)
  $S\leftarrow\{\}$
  for each sentence $x$ in $D$ do
     $H_{x}\leftarrow$ GetLastHiddenStates($x$, $\Theta$)
     $N_{x}\leftarrow$ CountAttentionMasked($x$)
     $R_{x}\leftarrow\frac{\textsc{Sum}(H_{x})}{N_{x}}$
     $S\leftarrow S\cup\{R_{x}\}$
  end for
  $W,b\leftarrow$ InitializeParameters()
  for each training epoch do
     $P_{U}\leftarrow(W\times S)+b$
     $loss\leftarrow$ ComputeLoss($P_{U}$, $F_{U}$)
     $W,b\leftarrow$ UpdateParametersWithConstraint($W,b,loss$)
  end for
  $A\leftarrow\{\}$
  for each sentence representation $s$ in $S$ do
     $A_{m}\leftarrow(W\times s)+b$
     $A\leftarrow A\cup\{A_{m}\}$
  end for
Appendix C Language Family Experiment
Tables 7 and 8 provide a comprehensive analysis of language family performance across different training groups. These tables compare the accuracy percentages of the Multilingual BERT and XLM-RoBERTa models, respectively. The results displayed in the tables elucidate the models’ capabilities in generalizing from the training data to unseen languages. A clear trend that can be observed is the improvement in performance as the training groups progress from 1 to 4, which suggests that the models benefit from exposure to a wider variety of language families during training. The ’Average’ row at the bottom of each table indicates the mean accuracy across all language families, providing an insight into the overall performance enhancement achieved by each model with incremental training diversity.
Language Family Train Group 1 Train Group 2 Train Group 3 Train Group 4
Afro-Asiatic 52.82% 52.93% 61.26% 61.00%
Atlantic-Congo 65.71% 68.08% 70.62% 71.79%
Austroasiatic 64.77% 66.78% 69.72% 70.16%
Austronesian 66.88% 68.66% 72.06% 72.19%
Dravidian 64.74% 67.97% 70.93% 71.41%
Indo-European 67.50% 68.61% 72.53% 72.95%
Japonic 72.11% 71.98% 75.80% 75.67%
Kartvelian 68.91% 68.89% 72.46% 72.32%
Koreanic 64.80% 66.46% 70.04% 69.91%
Mongolic-Khitan 63.11% 66.44% 69.71% 69.59%
Sino-Tibetan 62.65% 66.29% 68.79% 70.33%
Tai-Kadai 63.52% 67.89% 70.23% 71.34%
Turkic 54.69% 56.91% 63.54% 64.05%
Uralic 71.49% 71.27% 75.33% 75.15%
Average 65.54% 67.07% 71.04% 71.43%
Table 7: Multilingual BERT Performance of Language Families Across Training Groups
Language Family Train Group 1 Train Group 2 Train Group 3 Train Group 4
Afro-Asiatic 75.74% 76.23% 85.56% 85.39%
Atlantic-Congo 70.86% 72.38% 83.24% 82.73%
Austroasiatic 74.85% 76.04% 83.91% 83.59%
Austronesian 78.94% 79.83% 84.77% 84.69%
Dravidian 81.49% 82.20% 85.41% 85.43%
Indo-European 80.31% 81.21% 83.26% 83.47%
Japonic 80.21% 81.36% 82.67% 83.15%
Kartvelian 80.40% 81.53% 82.79% 83.27%
Koreanic 79.74% 80.91% 82.14% 82.61%
Mongolic-Khitan 79.54% 81.00% 82.20% 82.65%
Sino-Tibetan 79.25% 81.00% 82.14% 82.58%
Tai-Kadai 79.08% 80.81% 81.90% 82.35%
Turkic 79.20% 80.90% 81.96% 82.39%
Uralic 79.24% 80.91% 81.92% 82.47%
Average 79.45% 80.48% 83.44% 83.62%
Table 8: XLM-RoBERTa Performance of Language Families Across Training Groups